OpenStudy (speedsterpro):

Taylor complete the problem below. Describe and correct the error in the problem 2^-4*-2*-2*-2*-2=16 A. Taylor changed the negative exponent to a negative base. The problem should be 2^-4=1/2^4=1/2*2*2*2=1/16 B. Taylor didn't multiply correctly. The answer should be 16 C. Taylor made the answer positive. It should be -16 D. Taylor didn't multiply by -1. The problem should be 2^-4=(-1)1/2^4=(-1)1/2*2*2*2=1/16

OpenStudy (google):

Hm I don't understand this one

OpenStudy (speedsterpro):

Me either that is why I asked for your help lol

OpenStudy (google):

lol yeah sorry @sweetburger

OpenStudy (google):

@mathstudent55

OpenStudy (mathstudent55):

I don't understand what this means 2^-4*-2*-2*-2*-2=16

OpenStudy (mathstudent55):

Can you take a pic and post it? If not, can you use the equation editor? If not, can you draw it showing the exponents properly?

OpenStudy (speedsterpro):

First it is 2^-4

OpenStudy (speedsterpro):

But you have to correct it

OpenStudy (speedsterpro):

2*2*2*2=1/16

OpenStudy (mathstudent55):

Did you put a - where it should be =?

OpenStudy (mathstudent55):

The problem reads: \(2^{-4} = (-2) \times (-2) \times (-2) \times (-2) =16\) The mistake is that a negative exponent does not mean to make the base negative. A negative exponent means to set up a fraction. \(\Large a^{-n} = \dfrac{1}{a^n} \) \(2^{-4} = \dfrac{1}{2^4} = \dfrac{1}{2 \times 2 \times 2 \times 2} = \dfrac{1}{16}\)

OpenStudy (mathstudent55):

Now pick the option that is the same as the last calculation above.

OpenStudy (speedsterpro):

So which one would be the correct answer